OpenStudy (anonymous):

\[F(x, y, z) = {GmM \over {\sqrt{x^2+y^2+z^2}}}\] calculate \[\frac{\partial^2 F}{\partial x^2} + \frac{\partial^2 F}{\partial y^2} + \frac{\partial^2 F}{\partial z^2}\]

OpenStudy (dumbcow):

do you mean partial derivatives with respect to x^2, y^2, z^2 ?

OpenStudy (dumbcow):

or is it just 2nd derivatives with respect to x,y,z

OpenStudy (anonymous):

yes, I think that notation means take the second derivatives

OpenStudy (anonymous):

= 2 ( x² + y² + z²) / √(x² + y² + z²)

OpenStudy (anonymous):

I found it 0

OpenStudy (anonymous):

it is odd

OpenStudy (anonymous):

what do you mean by odd?

OpenStudy (anonymous):

strange..

OpenStudy (anonymous):

but it coul be thoug..

OpenStudy (anonymous):

\[\frac{\partial F}{\partial x} = -\frac{xGmM}{(x^2+y^2+z^2)^{3/2}}\\\frac{\partial^2F}{\partial x^2} = \frac{GmM(2x^2-y^2-z^2)}{(x^2+y^2+z^2)^{5/2}}\] does this look right?

OpenStudy (anonymous):

you are right..

OpenStudy (anonymous):

and total is zero..

OpenStudy (anonymous):

It says to simplify the answer as much as possible, I guess that 0 will work :)

OpenStudy (anonymous):

I found \[\frac{\partial^2 F}{\partial x^2}=3x^2(x^2+y^+z^2)^{-5/2}-(x^2+y^2+z^2){-3/2}\] it is the same thing.

OpenStudy (anonymous):

your answer is big probably zero..